# Design of a Wearable Parallel Electrical Impedance Imaging System for Healthcare

### v1

@@ 3.1 System Performance Evaluation
To verify the performance of the impedance acquisition device, the collected data was initially reviewed. In a water tank with a bottom diameter of 20 cm, 3 cm of water with a conductivity of 150 $\mu\text{S}/\text{cm}$ was added, and no objects were placed inside the tank. The impedance acquisition device was used to collect impedance data under two measurement modes: four-terminal serial and four-terminal semi-parallel, with a frequency range set from 8 kHz to 100 kHz. To avoid frequency leakage, when using a 16 MHz external crystal oscillator, the frequency was set to integer multiples of 1 MHz/1024. Fig.15 shows the data collected at 9766 Hz and 97656 Hz under the two measurement modes.
From the data acquisition results, it can be observed that the data quality was high across both measurement modes, which preliminarily verifies that the device supports a working frequency up to 97656 Hz.
@@
Subsequently, under the same experimental conditions, the system’s signal-to-noise ratio (SNR) and relative standard deviation (RSD) were measured in the water tank. For each frequency, 100 measurements were taken to calculate the SNR and RSD. The SNR was calculated using Eq.23 [25].
@@
Where $Z_{i}$ represents the value of each measurement, $\bar{Z}$ is their average, and N is the total number of measurements (in this case, N = 100). The $\sum_{i=1}^{N}Z_{i}^{2}$ in the formula represents the total signal strength, and $\sum_{i=1}^{N}(Z_{i}-\bar{Z})^{2}$ represents the noise component. The RSD was calculated using Eq.24 [25].
@@
Where $Z_{i}$ represents the value of each measurement, $\bar{Z}$ is their average, and N is the total number of measurements (again, N = 100). The standard deviation $\sqrt{\frac{1}{N}\sum_{i=1}^{N}(Z_{i}-\bar{Z})^{2}}$ of the measurement data, divided by the mean $\bar{Z}$, indicates the degree of dispersion or variability of the data.
Fig.16(a) and Fig.16(b) show the SNR and RSD at various frequencies, respectively. In both the four-terminal serial and four-terminal semi-parallel measurement modes, the SNR values remain between 50 dB and 70 dB, while the RSD values stay below 0.4% across all frequencies. This indicates that the measurement data exhibit small fluctuations and high stability.
@@
Finally, under the same experimental conditions, the system’s reciprocity error (RE) was measured in the water tank using adjacent excitation and adjacent measurement modes. The RE was calculated using Eq.25 [28].
@@
Where $v(c,d)\mid I(a,b)$ represents the voltage difference measured between electrodes c and d when current is applied between electrodes a and b, and $v(a,b)\mid I(c,d)$ represents the voltage difference measured between electrodes a and b when current is applied between electrodes c and d.
Fig.17 shows the RE at various frequencies. At all frequencies, the RE values are below 2%, indicating that the system exhibits good reciprocity.
@@ 3.2 Water Tank Phantom Imaging Results
This study employs the difference imaging method, which requires two frames of data: one as reference data and the other as target data. Difference imaging consists of time-difference imaging and frequency-difference imaging. Time-difference imaging is based on impedance data collected at different time points, while frequency-difference imaging is based on impedance data collected at different frequency points. In subsequent experiments, sweet potatoes and plastic bottles of various shapes and sizes were placed in water, and impedance magnitude data collected by the impedance acquisition device were used for time-difference imaging experiments. Image reconstruction was carried out using the one-step Gauss-Newton method and implemented in the EIDORS toolbox of MATLAB R2020b (MathWorks, USA). The experimental results demonstrate the imaging performance of the system in time-difference amplitude imaging mode under the two-terminal serial, four-terminal serial, and four-terminal semi-parallel measurement modes, and compare the performance across different frequencies. The specific imaging results are shown in Fig.18.
Figure 18: Comparison of Time-Difference Amplitude Images across Modes and Frequencies.
@@
In the imaging map, the blue areas represent high impedance, while the red areas represent low impedance. According to the experimental results, the imaging color of the sweet potato changes with the test frequency: at lower frequencies, the sweet potato appears blue, while at higher frequencies, it appears red. This is because the impedance of biological tissues decreases as the frequency increases. At low frequencies, the impedance of the sweet potato is higher than that of water, while at higher frequencies, it is lower than that of water. Furthermore, the impedance difference between the sweet potato and water is more pronounced at both low and high frequencies, resulting in clearer imaging. In contrast, the plastic bottle in the tank always appears blue and remains clear across all frequencies, as its impedance does not change with frequency and is consistently higher than that of water.
In subsequent experiments, frequency-difference imaging was performed using impedance magnitude data collected by the impedance acquisition device, with the sweet potato submerged in water. The same device was then used to collect data for frequency-difference imaging tests with submerged potatoes and plastic bottles. The experimental results demonstrate the system’s imaging performance in frequency-difference amplitude imaging mode under the two-terminal serial, four-terminal serial, and four-terminal semi-parallel measurement modes, and compare the performance of reference data and target data at different frequencies. The specific imaging results are shown in Fig.19.
Figure 19: Comparison of Frequency-Difference Amplitude Images across Modes and Frequencies.
@@
The imaging results show that when a sweet potato is placed in water for frequency-difference imaging, the region corresponding to it appears red, indicating a lower impedance value. This is because the impedance of the sweet potato decreases with increasing frequency. Similarly, when a potato and plastic bottle are placed in water for frequency-difference imaging, the region corresponding to the potato also appears red, while the region corresponding to the plastic bottle shows no color. This is because the impedance of the potato decreases with increasing frequency, while the impedance of the plastic bottle remains unchanged and unaffected by frequency. Furthermore, when the reference data frequency is lower and the target data frequency is higher (i.e., when the frequency difference is larger), the imaging of the sweet potato and potato is clearer. This is because a larger frequency difference results in a more significant impedance variation in the sweet potato and potato. Moreover, compared to the two-terminal measurement method, the four-terminal method more accurately reflects the size and shape of the sweet potato. This is because the four-terminal method includes more impedance information in each data frame and effectively suppresses contact impedance interference.
Next, time-difference imaging experiments using the imaginary component of impedance data were conducted with the submerged sweet potato, followed by similar experiments with the submerged plastic bottle. The experimental results demonstrate the imaging performance of the device in time-difference imaginary part imaging mode under the two-terminal serial, four-terminal serial, and four-terminal semi-parallel measurement modes, and compare the device’s performance at different frequencies. The specific imaging results are shown in Fig.20.
Figure 20: Comparison of Time-Difference Imaginary-Part Images across Modes and Frequencies.
The imaging results show that the region corresponding to the carrot appears red, while the region corresponding to the plastic bottle appears blue. This is because the carrot, as a biological entity made up of cells, exhibits capacitive characteristics, while the plastic bottle, as a non-biological material, shows different impedance characteristics. The significance of imaginary part imaging lies in its ability to reveal the phase information of impedance data, in contrast to amplitude imaging. In the biomedical field, due to differences in the conductivity and capacitive properties of biological tissues and other materials, imaginary part imaging provides additional spatial information, which helps distinguish between different types of tissues or materials.
@@ 3.3 Calf Cross-Section Imaging Results
In the experiment to reconstruct the cross-sectional impedance image of human calf tissue, data collection was conducted on a 28-year-old male subject with a Body Mass Index (BMI) of 20.5. Sixteen copper electrodes were evenly arranged around the calf using a PCB electrode band and connected to the impedance acquisition device, as shown in the Fig.21(a). The impedance acquisition device was then used to collect the impedance magnitude data of the calf tissue at two frequencies, 8789 Hz and 97656 Hz, under the four-terminal serial measurement mode. Finally, frequency-difference imaging was applied to reconstruct the cross-sectional image of the calf. Image reconstruction was performed using the one-step Gauss-Newton method, implemented in the EIDORS toolbox of MATLAB R2020b (MathWorks, USA). Fig.22 shows the comparison of the impedance frequency-difference imaging results of the calf cross-section with the anatomical structures.
The experimental results show that frequency-difference imaging can effectively reflect the tissue structure of the calf cross-section. For example, the tibia and fibula regions appear blue, while the blood vessels and nerve regions appear red. This is because different tissues (such as muscle, fat, blood, and bone) exhibit varying conductivity changes at different frequencies. Blood vessels and nerves, which contain more water and electrolytes, show significant impedance variation with frequency, while the impedance of the fibula and tibia changes less with frequency. These results demonstrate the broad potential of EIT for non-invasive detection and tissue identification.
@@ 3.4 Pulmonary Respiration Imaging Results
In the experiment of reconstructing dynamic lung breath images, data collection was similarly conducted on a 28-year-old male subject, with a measurement frequency of 97656 Hz. Sixteen copper electrodes were evenly arranged around the thorax using a PCB electrode band and connected to the impedance acquisition device, as shown in the Fig.21(b). Next, the subject was instructed to hold their breath in a resting state, and the device collected impedance data as reference data. The subject then resumed normal breathing, and the device continued to collect impedance data as target data. Finally, time-difference imaging was used to reconstruct the dynamic changes in lung breathing. The reconstruction process used one-step Gauss-Newton method. Fig.23 and Fig.24 show the imaging results of lung breathing changes under four-terminal serial and four-terminal semi-parallel measurement modes.
The imaging maps include five complete breathing cycles, with each row representing one full breathing cycle and each column displaying images from different stages of breathing. In the imaging maps, the blue areas represent high-impedance regions, while the red areas represent low-impedance regions. From these images, it is clear that as breathing progresses, the impedance values in different lung regions change. During the inhalation phase, gas enters the chest, and the impedance of the corresponding lung region increases, appearing as blue. During the exhalation phase, gas exits the chest, and the impedance of the corresponding lung region decreases, appearing as red. The results indicate that the impedance acquisition device has excellent accuracy and stability, and is capable of effectively capturing the dynamic changes in lung breathing.
@@
Figure 24: Pulmonary Respiration Imaging with Four-Terminal Semi-Parallel Mode.
@@ 4 Conclusion
To address the issues of high cost and bulkiness in existing electrical impedance tomography (EIT) devices, this study presents a high-precision, high-speed, low-cost, and portable impedance acquisition device. The main research contents include:
@@
    Development of a graphical user interface (GUI) using PyQt5, streamlining user operations and facilitating real-time imaging and data acquisition.
@@
    Performance evaluation of the impedance acquisition device. Initial experiments using a water tank phantom demonstrated effective imaging capabilities across various measurement modes (two-terminal and four-terminal, serial and parallel) and imaging modes (time-difference and frequency-difference) up to 100 kHz. Subsequent tests involved successful dynamic lung respiration imaging using the time-difference method, effectively capturing respiratory movements. Finally, frequency-difference imaging was utilized for cross-sectional imaging of calf tissue, successfully distinguishing anatomical features such as bones, blood vessels, and nerves.

### v2

@@ 3.1 System Performance Evaluation
To evaluate the performance of the impedance acquisition device, a preliminary inspection of the collected data was conducted. A water tank with a bottom diameter of 20 cm was filled with water to a depth of approximately 3 cm. The water had a conductivity of 150 $\mu\text{S}/\text{cm}$, and no objects were placed in the tank during the measurement. Impedance data were collected using the impedance acquisition device under two measurement modes: four-terminal serial and four-terminal parallel. The frequency range was set from 8 kHz to 100 kHz, and to minimize frequency leakage, the excitation frequencies within this range were configured as integer multiples of 1 MHz/1024 when using a 16 MHz crystal oscillator. Fig.15 presents the data collected at frequencies of 9766 Hz and 97656 Hz under both measurement modes.
From the data acquisition results, both measurement modes demonstrated high data standardization, providing preliminary validation that the device can operate reliably at frequencies up to 97656 Hz.
@@
Subsequently, under the same experimental conditions, the system’s signal-to-noise ratio (SNR) and relative standard deviation (RSD) were measured in the water tank. At each frequency, 100 measurements were conducted to calculate the SNR and RSD. The SNR was calculated using Eq.14, while the RSD was determined according to Eq.15 [25].
@@
Here, $Z_{i}$ denotes the value of each individual measurement, $\bar{Z}$ represents the average of these values, and N is the number of measurements (in this case, N = 100).
Fig.16(a) and Fig.16(b) show the SNR and RSD measured at various frequencies, respectively. In both the four-terminal serial and four-terminal parallel measurement modes, the SNR remains within the range of 50 to 70 dB, while the RSD stays below 0.3% across all frequencies. These results indicate minimal fluctuation and high stability in the measurement data.
@@
Finally, under the same experimental conditions, the system’s reciprocity error (RE) was measured in the water tank using the adjacent excitation and adjacent measurement modes. The RE was calculated according to Eq.16 [28].
@@
Here, $v(c,d)\mid I(a,b)$ denotes the voltage difference measured between electrodes c and d when current is applied between electrodes a and b, while $v(a,b)\mid I(c,d)$ denotes the voltage difference measured between electrodes a and b when current is applied between electrodes c and d.
Fig.17 shows the RE measured at various frequencies. At all frequencies, the RE remains below 0.8%, demonstrating that the system possesses good reciprocity.
@@ 3.2 Water Tank Phantom Imaging Results
The difference imaging method requires two frames of data: one serving as the reference and the other as the target. There are two types of difference imaging: time-difference imaging and frequency-difference imaging. Time-difference imaging utilizes impedance data acquired at different time points, whereas frequency-difference imaging uses data collected at different frequencies.
@@
In subsequent experiments, impedance magnitude data collected by the impedance acquisition device were used to perform time-difference imaging on sweet potatoes and plastic bottles of various shapes and sizes submerged in water. Fig.18 presents the imaging results. The imaging reconstruction was performed using the one-step Gauss-Newton method, implemented with the EIDORS toolbox in MATLAB R2020b (MathWorks, USA).
@@
According to the experimental results, the imaging color of the sweet potato changes with the test frequency: it appears blue at lower frequencies and turns red at higher frequencies. This is due to the frequency-dependent impedance of biological tissues, which decreases as the frequency increases. At low frequencies, the impedance of the sweet potato is higher than that of water, whereas at high frequencies, it becomes lower. Moreover, the impedance contrast between the sweet potato and water is more pronounced at both low and high frequencies, leading to clearer imaging results. In contrast, the plastic bottle in the water tank consistently appears blue across all frequencies, as its impedance remains constant and is significantly higher than that of water.
Additionally, impedance magnitude data from the same acquisition device were employed for frequency-difference imaging, initially on sweet potatoes submerged in water and subsequently on potatoes and plastic bottles. Fig.19 illustrates the outcome of the imaging.
@@
The imaging results demonstrate that when a sweet potato is placed in water for frequency-difference imaging, its corresponding region appears red, indicating a lower impedance value. This occurs because the impedance of the sweet potato decreases as the frequency increases. Similarly, when a potato and a plastic bottle are placed in the water, the region corresponding to the potato also appears red, whereas the plastic bottle’s region remains uncolored. This is due to the fact that the potato’s impedance likewise decreases with increasing frequency, while the impedance of the plastic bottle remains constant and is unaffected by frequency changes.
Finally, using the imaginary part of impedance data gathered by the same device, time-difference imaging was conducted on submerged carrots, followed by similar imaging on plastic bottles. Fig.20 demonstrates the imaging results.
The imaging results indicate that the region corresponding to the carrot appears red, while the area corresponding to the plastic bottle appears blue. This distinction arises because the carrot, as a biological tissue composed of cells, exhibits capacitive properties, whereas the plastic bottle, being a non-biological material, displays distinct electrical impedance characteristics.
Figure 18: Time-Difference Imaging Result Based on Impedance Amplitude Data.
Figure 19: Frequency-Difference Imaging Result Based on Impedance Amplitude Data.
Figure 20: Time-Difference Imaging Result Based on Imaginary Part of Impedance Data.
@@ 3.3 Calf Cross-Section Imaging Results
In the experiment to reconstruct the cross-sectional impedance image of human calf tissue, data collection was conducted on a 28-year-old male subject with a Body Mass Index (BMI) of 20.5. Firstly, Sixteen copper electrodes were evenly arranged around the calf using a PCB-based electrode band and connected to the impedance acquisition device, as shown in Fig.21(a). The device then measured the impedance magnitude of the calf tissue at two frequencies—8,789 Hz and 97,656 Hz—using the four-terminal serial measurement mode. Finally, frequency-difference imaging was employed to reconstruct the cross-sectional image of the calf. The reconstruction was carried out using the one-step Gauss-Newton method, implemented in the EIDORS toolbox within MATLAB R2020b (MathWorks, USA). Fig.22 presents a comparison between the impedance frequency-difference imaging results and the anatomical structures of the calf cross-section.
The experimental results demonstrate that frequency-difference imaging can effectively reveal the tissue structure of the calf cross-section. For example, the tibia and fibula regions appear blue, while blood vessels and nerve regions appear red. This contrast arises because blood vessels and nerves contain higher levels of water and electrolytes, resulting in more pronounced impedance variations with frequency. In contrast, the impedance of the tibia and fibula remains relatively stable across different frequencies.
@@ 3.4 Pulmonary Respiration Imaging Results
In the experiment of reconstructing dynamic lung breath images, data collection was similarly conducted on a 28-year-old male subject, with a measurement frequency of 97656 Hz. Firstly, Sixteen copper electrodes were evenly arranged around the thorax using a PCB-based electrode band and connected to the impedance acquisition device, as shown in Fig.21(b). The subject was first instructed to hold their breath in a resting state, during which impedance data were collected as reference data. The subject then resumed normal breathing, and the device continued recording impedance data as target data. Finally, time-difference imaging was used to reconstruct the dynamic changes in lung respiration. The reconstruction was performed using the one-step Gauss-Newton method. Fig.23 and Fig.24 show the imaging results of pulmonary ventilation changes under the four-terminal serial and four-terminal parallel measurement modes.
The imaging maps display five complete breathing cycles, with each row corresponding to a full cycle and each column representing different phases of respiration. These images clearly illustrate the dynamic changes in impedance across various lung regions throughout the respiratory process. During inhalation, as air enters the thoracic cavity, the impedance in the relevant lung areas increases, appearing in blue. Conversely, during exhalation, as air leaves the thoracic cavity, the impedance decreases, which is visualized in red.
@@
Figure 24: Pulmonary Respiration Imaging with Four-Terminal Parallel Mode.
@@ 4 Conclusion
To address the limitations of existing EIT devices, such as high cost and bulky design, this study proposes a portable impedance acquisition device characterized by high precision, fast measurement speed, and affordability. The primary research areas include:
@@
    Developed a user-friendly GUI with PyQt5 to simplify operation procedures and support efficient real-time imaging and data collection.
@@
    Systematic performance evaluations were conducted on the EIT device, demonstrating excellent results: the SNR exceeded 50 dB, the RSD was below 0.3%, and the RE was less than 0.8%. To further assess its imaging capabilities, three experiments were carried out. First, a water tank phantom experiment showed that the device provided stable and accurate imaging at frequencies below 100 kHz. Second, time-difference imaging was used to visualize lung ventilation, successfully capturing the dynamic changes associated with the respiratory cycle. Third, frequency-difference imaging was applied to cross-sectional imaging of the calf, with the results indicating the preliminary differentiation of anatomical structures such as bones, blood vessels, and nerves.